OpenStudy (anonymous):

Solve for x by changing to exponential form. Round your answer to three decimal places. 9Ln x= 13 .

OpenStudy (anonymous):

By reviewing logrithmic properties, you will see that you can put the 9 coefficient as the expontent to the bas of the ln, to make x^2. Then if you exponetiate both sides of the equation you will be able to cancel out the e and ln on the left side of the equation. In other words e^ln(x^2) = e^13... which simlifies to x^2 = e^13.

OpenStudy (anonymous):

sorry, that should be x^9, not x^2

OpenStudy (anonymous):

yeah but from there am stuck i need to find an answer for it:/

OpenStudy (anonymous):

the next step would be to solve for x... from there you would probably need a calculator to find your decimal answer... e equals about 2.71, but it is better if you use the e function on your calculator to get more accurate results

OpenStudy (anonymous):

\[x^9 = e^{13}\]\[x = \sqrt[9]{e^{13}} = e^{13 / 9}\] From there I would punch it in my calculator.

OpenStudy (anonymous):

thanks! i didnt know how to do that last step

OpenStudy (anonymous):

answer :4.239

OpenStudy (anonymous):

awesome!